# Power law $\alpha$-Starobinsky inflation

### v1

@@ Abstract
In this work we consider a generalization of Starobinsky inflation obtained by combining power law ($R^{\beta}$), and $\alpha$-Starobinsky inflation ($E$-model). The Einstein frame potential for this model is that of power law Starobinsky inflation modified by a parameter $\alpha$ in the exponential. After computing power spectra for scalar and tensor perturbations numerically, we perform MCMC analysis to put constraints on the potential parameter $\alpha$, $\beta$ and $M$, and the number of e-foldings $N_{pivot}$ during inflation, using Planck-2018, BICEP/Keck (BK18) and other LSS observations. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$, $\beta=1.969^{+0.020}_{-0.023}$, $M=\left(3.54^{+2.62}_{-1.73}\right)\times 10^{-5}$ and $N_{pivot}=47\pm{10}$. We compute the Bayesian evidences for our proposed model, power law Starobinsky inflation, $\alpha$-Starobinsky inflation and Starobinsky inflation. Considering the Starobinsky model as the base model, we calculate the Bayes factor and find that our proposed model is preferred by the CMB and LSS observations.
@@ I INTRODUCTION
While inflation is generally seen as a successful idea, we still have not found a unique model of inflation that fits well within observational limits which have become more precise overtime. Since its invent, many different inflationary models have been proposed and examined Martin:2006rs ,Martin:2013tda with observational data. Analysis of recent Planck 2018 results Planck:2018jri shows that the simplest classic inflaton potentials, $\frac{1}{2}m^{2}\phi^{2}$ and $\lambda\phi^{4}$, are strongly ruled out. As proposed by Starobinsky Starobinsky:1980te , inflation can also be achieved without scalar field by adding $\frac{1}{M}R^{2}$ term, $R$ being the Ricci scalar, in the Einstein-Hilbert action. Named as Starobinsky inflation, this model is in remarkable agreement with Planck-2018 Planck:2018jri and BICEP/Keck BICEP:2021xfz ; Tristram:2021tvh observations as it predicts lower values for tensor-to-scalar ratio $r$. Originally described in the Jordan frame, Starobinsky inflation can be reinterpreted in the Einstein frame through a conformal transformation, where it looks like a single field inflationary model with a potential that flattens out at large field values ensuring a slow and sustained period of inflation. Interestingly this potential in the Einstein frame can be easily derived from no-scale supergravity with a non-compact $SU(2,1)/SU(2)\times U(1)$ symmetry Ellis:2013xoa , where we have a modulus field fixed by the other dynamics and the inflaton field is a part of the chiral superfield with a minimal Wess-Zumino superpotential.
@@
Several generalizations of Starobinsky inflation have been proposed, such as, power law Starobinsky inflation having $R^{\beta}$ term in the Einstein-Hilbert action Schmidt:1989zz ; Maeda:1988ab ; Muller:1989rp ; Gottlober:1992rg ; DeFelice:2010aj ; Nojiri:2010wj ; Nojiri:2017ncd , which is based on higher-order metric theories of gravity, and $\alpha$-Starobinsky inflation ($E$-model) based on supergravity inspired deformation of the inflaton potential in the Einstein frame Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2013yoa . Although inspired by higher-order metric theories of gravity, the potential for the power law Starobinsky inflation can also be derived from no-scale supergravity Chakravarty:2014yda . Power law Starobinsky inflation gained traction in 2014 when BICEP2 reported a large value of tensor-to-scalar ratio $r=0.2^{+0.07}_{-0.05}$ BICEP2:2014owc , as it was shown that this model generates a large $r$ compared to Starobinsky inflation for $\beta$ slightly less than $2$ Codello:2014sua ; Costa:2014lta ; Martin:2014lra . The primordial origin of BICEP2 signal of $B$-mode polarization was ruled out later and it was found the it is due to the foreground dust emission Planck:2014dmk . Observational constraints on $R^{\beta}$ inflation have been studied in Chakravarty:2014yda ; Motohashi:2014tra ; Odintsov:2022bpg ; Meza:2021xuq ; Saini:2023uqc and it is found that the current data allows a slight deviation from $\beta=2$ (see Saini:2023uqc for a detailed statistical analysis with Planck-2018 and BICEP3/Keck array data). In $\alpha$-Starobinsky inflation the scalaron potential in the Einstein frame contains a parameter $\alpha$ in the exponential that modifies the predictions for $n_{s}$ and $r$. The potential for this model can be obtained by generalizing the coefficient of the logarithm of the volume modulus field in the Kähler potential in $SU(2,1)/SU(2)\times U(1)$ no-scale supergravity with a suitable choice of the superpotential having both the volume modulus field and the chiral superfield. The parameter $\alpha$ is these models is related to the geometry of the Kähler manifold. The observational constraints on $\alpha$-Starobinsky inflation from Planck and BICEP/Keck array data are studied in Planck:2015sxf ; Planck:2018jri ; Ueno:2016dim ; Ellis:2020xmk ; Ellis:2021kad ; ElBourakadi:2022lqf ; Sarkar:2021ird ; Saini:2024mun , and it is found that the current data is not sufficient to constrain the parameter $\alpha$. However, an upper limit on $\alpha$ i.e., $\alpha\leq 39.81$ with $95\%$ C. L. is obtained in Saini:2024mun from Planck-2018 and BICEP/Keck array observations by performing a detailed statistical analysis. A two field inflation model with both the field having $\alpha$-attractor potentials have also been studied in Rodrigues:2020fle . The $E$-model potential in brane inflation has also been studied in Sabir:2019wel . In SantosdaCosta:2020dyl the variations from Starobinsky potential in the Einstein frame have been studied based on a potential derived from brane inflation.
@@ II Power law $\alpha$-Starobinsky Model
The modified Starobinsky model, which includes an $R^{\beta}$ correction, has been analyzed in Chakravarty:2014yda . The corresponding action in Jordan frame is given by,
$S_{J}=\frac{-M_{Pl}^{2}}{2}\int\sqrt{-g}f(R)d^{4}x,\quad\quad f(R)=\left(R+\frac{1}{6M^{2}}\frac{R^{\beta}}{M_{Pl}^{2\beta-2}}\right)$ (1)
@@
To obtain the action in the Einstein frame we use conformal transformations $\tilde{g}_{\mu\nu}(x)=\Omega(x)g_{\mu\nu}(x)$, where a tilde denotes the quantities in the Einstein frame and the conformal factor $\Omega(x)$ is defined as
$\Omega=F=\frac{\partial f(R)}{\partial R}.$ (2)
We also need to introduce a new scalar field $\chi$ defined in terms of the conformal factor as
@@
Using Eqns. (2) and (3) the action (1) can be transformed to the Einstein frame as
$S_{E}=\int d^{4}x\sqrt{-\tilde{g}}\left(\frac{-1}{2}\tilde{R}+\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\chi}\partial_{\nu}{\chi}+V(\chi)\right),$ (4)
@@
$V(\chi)=\frac{(RF(R)-f(R))}{2F(R)^{2}},$ (5)
Thus, the potential for the power law $\alpha$ Starobinsky Model can be expressed as
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right]\times\left(1-\exp\left(-\frac{2\chi}{\sqrt{6\alpha}}\right)\right)^{\frac{\beta}{\beta-1}}.\$ (6)
In the large field limit $\chi\gg 2\sqrt{6}$ and $21<\beta<$ 2 the potential from Eq. (6) simplifies to
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right]$ (7)
@@ II.1 No-scale supergravity realization of the power law $\alpha$-Starobinsky Model
The Einstein frame potential for the generalized power law $\frac{1}{M^{2}}R^{\beta}$ model has been obtained in Chakravarty:2014yda by incorporating a $(\Phi+\Phi^{*})^{n}$ term into the minimal no-scale Kähler potenial along with a Wess Zumino superpotential. The potential for power law $\alpha$-Starobinsky model can be obtained by generalizing the coefficient of the volume modulus field in the Kähler potential and following the same approach as in Chakravarty:2014yda . So the Kähler potenial with parameter $\alpha$ is given as
$K=-3\alpha\ln\left[T+T^{*}-\frac{1}{2}(\phi+\phi^{*})^{n}\right]$ (8)
@@
$\mathcal{L}_{E}=\frac{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}{4\left[c-\frac{(2\phi)^{n}}{12}\right]^{2}}(\partial_{\mu}\phi)^{2}-\frac{4(2\phi)^{2-n}}{\alpha n(n-1)\left[c-\frac{(2\phi)^{n}}{12}\right]^{3\alpha-1}}\left|\frac{\partial W}{\partial\phi}\right|^{2}$ (9)
@@
$\frac{\partial\chi}{\partial\phi}=\frac{\sqrt{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}}{2\left(c-\frac{(2\phi)^{n}}{12}\right)}$ (10)
By integrating above equation:
$\phi\approx\frac{1}{2}\exp\left(\frac{2\chi}{\sqrt{3n\alpha}}\right)\left[1+\frac{6c(n+1)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right]$ (11)
Substituting above equation and Wess-Zumino superpotential $W=\frac{\mu}{2}\phi^{2}-\frac{\lambda}{3}\phi^{3}$ in the potential term of Lagrangian $\mathcal{L}_{E}$, we get
@@
×$\quad\times\left[1-\frac{2\mu}{\lambda}\exp\left(-\frac{2\chi}{\sqrt{3n\alpha}}\right)-\frac{9c(n^{2}-n-2)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right]^{2}$ (12)
we have assumed $21<n<$ 2, and in the large field limit $\chi\gg\sqrt{3n}$ above equation is equivalent to
$V=\frac{(12)^{3\alpha-1}\mu^{2}}{\alpha n(n-1)(-1)^{3\alpha-1}}\exp\left[\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{3\sqrt{2}(2-n\alpha)}{\sqrt{n}}\right)\right]$ (13)
By comparing the power law potential 7 and SUGRA potential 13, we can show the relation between the parameters of the two models.
@@
  • For $\beta<2$ (e.g.,$\alpha=20$, $\beta=1.90$),the potential becomes steeper. This may correspond to scenarios with larger tensor to scalar ratio in comparison to $\beta=2$.
  • For $\beta>2$ (e.g.,$\alpha=10$, $\beta=2.05$), the potential initially rises, reaches a maximum, and then decreases sharply, transitioning away from the plateau into a steeper.
@@
  • Larger values of $\alpha$ (e.g.,$\alpha=20$) stretch the potential horizontally and amplify the steepness of the potential for a fixed $\beta$, causing potential to grow more rapidly with $\chi$.
  • Smaller values of $\alpha$ (e.g.,$\alpha=0.34$) result in more gradual increases in potential.

### v2

@@ Abstract
In this work we consider a generalization of Starobinsky inflation obtained by combining power law ($R^{\beta}$), and $\alpha$-Starobinsky inflation ($E$-model). The Einstein frame potential for this model is that of power law Starobinsky inflation modified by a parameter $\alpha$ in the exponential. After computing power spectra for scalar and tensor perturbations numerically, we perform MCMC analysis to put constraints on the potential parameter $\alpha$, $\beta$ and $M$, and the number of e-foldings $N_{pivot}$ during inflation, using Planck-2018, BICEP/Keck (BK18) and other LSS observations. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$, $\beta=1.969^{+0.020}_{-0.023}$, $M=\left(3.54^{+2.62}_{-1.73}\right)\times 10^{-5}$ and $N_{pivot}=47\pm{10}$. With these mean values of the potential parameters $\alpha$ and $\beta$, and varying $N_{pivot}$ between $40$ to $55$, we also find that the $r-n_{s}$ predictions of our model lie well within the $1\sigma$ bounds of joint constraints from combined analysis of ACT, Planck-2018, BICEP and BAO observations. We compute the Bayesian evidences for our proposed model, power law Starobinsky inflation, $\alpha$-Starobinsky inflation and Starobinsky inflation. Considering the Starobinsky model as the base model, we calculate the Bayes factor and find that our proposed model is preferred by the CMB and LSS observations.
@@ I INTRODUCTION
While inflation is generally seen as a successful idea, we still have not found a unique model of inflation that fits well within observational limits which have become more precise overtime. Since its invent, many different inflationary models have been proposed and examined Martin:2006rs ,Martin:2013tda with observational data. Analysis of recent Planck 2018 results Planck:2018jri shows that the simplest classic inflaton potentials, $\frac{1}{2}m^{2}\phi^{2}$ and $\lambda\phi^{4}$, are strongly ruled out. As proposed by Starobinsky Starobinsky:1980te , inflation can also be achieved without scalar field by adding $\frac{1}{M}R^{2}$ term, $R$ being the Ricci scalar, in the Einstein-Hilbert action. Named as Starobinsky inflation, this model is in remarkable agreement with Planck-2018 Planck:2018jri and BICEP/Keck BICEP:2021xfz ; Tristram:2021tvh observations as it predicts lower values for tensor-to-scalar ratio $r$. Originally described in the Jordan frame, Starobinsky inflation can be reinterpreted in the Einstein frame through a conformal transformation, where it looks like a single field inflationary model with a potential that flattens out at large field values ensuring a slow and sustained period of inflation. Interestingly this potential in the Einstein frame can be easily derived from no-scale supergravity with a non-compact $SU(2,1)/SU(2)\times U(1)$ symmetry Ellis:2013xoa , where we have a modulus field fixed by the other dynamics and the inflaton field is a part of the chiral superfield with a minimal Wess-Zumino superpotential. This enables incorporating inflation with various other aspects of particle physics phenomenology Ellis:2016ipm ; Ellis:2017jcp ; Ellis:2019opr . However, analysis of recently reported Atacama Cosmology Telescope data ACT:2025fju , in combination with Planck 2018, BAO and DESY Y1 data, indicates that the Starobinsky model is disfavored at the $2\sigma$ level ACT:2025tim .
@@
Several generalizations of Starobinsky inflation have been proposed, such as, power law Starobinsky inflation having $R^{\beta}$ term in the Einstein-Hilbert action Schmidt:1989zz ; Maeda:1988ab ; Muller:1989rp ; Gottlober:1992rg ; DeFelice:2010aj ; Nojiri:2010wj ; Nojiri:2017ncd , which is based on higher-order metric theories of gravity, and $\alpha$-Starobinsky inflation ($E$-model) based on supergravity inspired deformation of the inflaton potential in the Einstein frame Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2013yoa . Although inspired by higher-order metric theories of gravity, the potential for the power law Starobinsky inflation can also be derived from no-scale supergravity Chakravarty:2014yda . Power law Starobinsky inflation gained traction in 2014 when BICEP2 reported a large value of tensor-to-scalar ratio $r=0.2^{+0.07}_{-0.05}$ BICEP2:2014owc , as it was shown that this model generates a large $r$ compared to Starobinsky inflation for $\beta$ slightly less than $2$ Codello:2014sua ; Costa:2014lta ; Martin:2014lra . The primordial origin of BICEP2 signal of $B$-mode polarization was ruled out later and it was found the it is due to the foreground dust emission Planck:2014dmk . Observational constraints on $R^{\beta}$ inflation have been studied in Chakravarty:2014yda ; Motohashi:2014tra ; Odintsov:2022bpg ; Meza:2021xuq ; Saini:2023uqc and it is found that the current data allows a slight deviation from $\beta=2$ (see Saini:2023uqc for a detailed statistical analysis with Planck-2018 and BICEP3/Keck array data). In $\alpha$-Starobinsky inflation the scalaron potential in the Einstein frame contains a parameter $\alpha$ in the exponential that modifies the predictions for $n_{s}$ and $r$. The potential for this model can be obtained by generalizing the coefficient of the logarithm of the volume modulus field in the Kähler potential in $SU(2,1)/SU(2)\times U(1)$ no-scale supergravity with a suitable choice of the superpotential having both the volume modulus field and the chiral superfield. The parameter $\alpha$ is these models is related to the geometry of the Kähler manifold. The observational constraints on $\alpha$-Starobinsky inflation from Planck and BICEP/Keck array data are studied in Planck:2015sxf ; Planck:2018jri ; Ueno:2016dim ; Ellis:2020xmk ; Ellis:2021kad ; ElBourakadi:2022lqf ; Sarkar:2021ird ; Saini:2024mun , and it is found that the current data is not sufficient to constrain the parameter $\alpha$. However, an upper limit on $\alpha$ i.e., $\alpha\leq 39.81$ with $95\%$ C. L. is obtained in Saini:2024mun from Planck-2018 and BICEP/Keck array observations by performing a detailed statistical analysis. A two field inflation model with both the field having $\alpha$-attractor potentials have also been studied in Rodrigues:2020fle . The $E$-model potential in brane inflation has also been studied in Sabir:2019wel . $R^{2}$ attractors and $\alpha$-attractors inflationary models as a subclass of $f(R)$ gravity in the Jordan frame have been studied in Odintsov:2020thl (see also Odintsov:2023ypt for various phenomenological aspects of these models). In SantosdaCosta:2020dyl the variations from Starobinsky potential in the Einstein frame have been studied based on a potential derived from brane inflation. Extensions of Starobinsky model with $R^{3}$, $R^{4}$ and $R^{3/2}$, in addition to $R^{2}$ term, have also been studied in Ivanov:2021chn in the light of Planck 2018 observations. It is shown in Gialamas:2025ofz that the Starobinsky inflation with an additional $R^{3}$ term is in full agreement with recent ACT observations.
@@ II Power law $\alpha$-Starobinsky Model
The modified Starobinsky model, which includes an $R^{\beta}$ correction, has been analyzed in DeFelice:2010aj ; Chakravarty:2014yda ; Motohashi:2014tra . The corresponding action in Jordan frame is given by,
$S_{J}=\frac{-M_{Pl}^{2}}{2}\int\sqrt{-g}f(R)d^{4}x,\quad\quad f(R)=\left(R+\frac{1}{6M^{2}}\frac{R^{\beta}}{M_{Pl}^{2\beta-2}}\right),$ (1)
@@
To obtain the action in the Einstein frame we use conformal transformations $\tilde{g}_{\mu\nu}(x)=\Omega(x)g_{\mu\nu}(x)$, where a tilde denotes the quantities in the Einstein frame. The Ricci scalar $R$ defined in the Jordan frame can be expressed in terms of the Einstein frame Ricci scalar $\tilde{R}$ via the conformal transformation connecting the two frames as
$R=\Omega\left(\tilde{R}+3\widetilde{\Box}\omega-\frac{3}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\omega}\partial_{\nu}{\omega}\right),$ (2)
here $\omega\equiv\ln\Omega$, $\widetilde{\Box}\omega\equiv\frac{1}{\sqrt{-\tilde{g}}}\partial_{\mu}\left(\sqrt{-\tilde{g}}\tilde{g}^{\mu\nu}\partial_{\nu}{\omega}\right)$ and $\partial_{\mu}{\omega}=\frac{\partial\omega}{\partial\tilde{x}^{\mu}}$.
To obtain action in the Einstein frame we choose $\Omega=F=\frac{\partial f(R)}{\partial R}$ and also introduce a new scalar field $\chi$ defined as
$\chi\equiv\frac{\sqrt{6\alpha}}{2}\ln F.$ (3)
This leads to
@@
Here, the parameter $\alpha$ is a dimensionless parameter, and it is introduced to obtain the potential of $\alpha$-Starobinsky inflation in the Einstein frame. Now, using Eq. (2) and (4) along with $\sqrt{-g}=\Omega^{-2}\sqrt{-\tilde{g}}$, the action (1) can be transformed to the Einstein-Hilbert form as
$S_{E}=\int d^{4}x\sqrt{-\tilde{g}}\left(-\frac{1}{2}\tilde{R}+\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\chi}\partial_{\nu}{\chi}+V(\chi)\right),$ (5)
@@
$V(\chi)=\frac{(RF(R)-f(R))}{2F(R)^{2}}.$ (6)
Thus, the potential for the power law $\alpha$-Starobinsky Model can be expressed as
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right]\times\left(1-\exp\left(-\frac{2\chi}{\sqrt{6\alpha}}\right)\right)^{\frac{\beta}{\beta-1}}.\\$ (7)
We can see that this potential reduces to the Starobinsky ($R^{2}$) potential in the Einstein frame for $\beta=2$ and $\alpha=1$. In the large field limit $\chi\gg 2\sqrt{6}$ and $1<\beta<2$ the potential from Eq. (7) simplifies to
$V({\chi})=(\frac{{\beta}-1}{2})\left(\frac{6M^{2}}{{\beta}^{\beta}}\right)^{\frac{1}{\beta-1}}\exp\left[{\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{2-\beta}{\beta-1}\right)}\right].$ (8)
@@ II.1 No-scale supergravity realization of the power law $\alpha$-Starobinsky Model
The Einstein frame potential for the generalized power law $\frac{1}{M^{2}}R^{\beta}$ model has been obtained in Chakravarty:2014yda by incorporating a $(\Phi+\Phi^{*})^{n}$ term into the minimal no-scale Kähler potential along with a Wess-Zumino superpotential Ellis:2013xoa ; Chakravarty:2014yda ; Ellis:2020xmk . The potential for power law $\alpha$-Starobinsky model can be obtained by generalizing the coefficient of the volume modulus field in the Kähler potential and following the same approach as in Chakravarty:2014yda . So the Kähler potential with parameter $\alpha$ is given as
$K=-3\alpha\ln\left(T+T^{*}-\frac{1}{2}(\phi+\phi^{*})^{n}\right),$ (9)
@@
$\mathcal{L}_{E}=\frac{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}{4\left[c-\frac{(2\phi)^{n}}{12}\right]^{2}}(\partial_{\mu}\phi)^{2}-\frac{4(2\phi)^{2-n}}{\alpha n(n-1)\left[c-\frac{(2\phi)^{n}}{12}\right]^{3\alpha-1}}\left|\frac{\partial W}{\partial\phi}\right|^{2}.$ (10)
@@
$\frac{\partial\chi}{\partial\phi}=\frac{\sqrt{n\alpha(2\phi)^{n-2}\left[c(n-1)+\frac{(2\phi)^{n}}{12}\right]}}{2\left(c-\frac{(2\phi)^{n}}{12}\right)}.$ (11)
By integrating above equation, we get
$\phi\approx\frac{1}{2}\exp\left(\frac{2\chi}{\sqrt{3n\alpha}}\right)\left[1+\frac{6c(n+1)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right].$ (12)
Substituting above equation and Wess-Zumino superpotential $W=\frac{\mu}{2}\phi^{2}-\frac{\lambda}{3}\phi^{3}$ Ellis:2013xoa ; Chakravarty:2014yda ; Ellis:2020xmk in the potential term of Lagrangian $\mathcal{L}_{E}$, we get
@@
$\quad\times\left[1-\frac{2\mu}{\lambda}\exp\left(-\frac{2\chi}{\sqrt{3n\alpha}}\right)-\frac{9c(n^{2}-n-2)}{n}\exp\left(-\frac{2n\chi}{\sqrt{3n\alpha}}\right)\right]^{2}.$ (13)
We have assumed $1<n<2$, and in the large field limit $\chi\gg\sqrt{3n}$ above equation is equivalent to
$V=\frac{(12)^{3\alpha-1}\mu^{2}}{\alpha n(n-1)(-1)^{3\alpha-1}}\exp\left[\frac{2\chi}{\sqrt{6\alpha}}\left(\frac{3\sqrt{2}(2-n\alpha)}{\sqrt{n}}\right)\right].$ (14)
By comparing the power law potential 8 and SUGRA potential 14, we can show the relation between the parameters of the two models.
@@
[…]
@@
[…]